# HOPSE: Scalable Higher-Order Positional and Structural Encoder for Combinatorial Representations

### v1

@@ 4.1 Methodology
This section describes HOPSE, a method for enriching cell features in combinatorial complexes with higher-order positional and structural encodings derived from Hasse graph decompositions. HOPSE enables expressive higher-order modeling without relying on the computational overhead of HOMP, while retaining its representational power. The full pipeline is illustrated in Figure 2, with components referenced using the corresponding labels (A–F) throughout this section.
Figure 2: HOPSE pipeline. (A) Considering the collection of neighborhoods ${\mathcal{N}}_{C}=\{{\mathcal{A}}_{0,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{0\to 2}\}$, the input combinatorial complex $\mathcal{T}$ is decomposed into the corresponding multi-attributed Hasse graphs expansions $\{\mathcal{J}_{{\mathcal{N}}}\}_{{\mathcal{N}}_{C}}$ –each one with $k$ associated signals. (B) Positional and structural encodings are computed for each Hasse graph, $\mathbf{G}(\mathcal{J}_{\mathcal{N}_{i}})$, as defined in Equation 1. (C) Higher-order PS encodings are derived via the target rank-aware aggregation $\psi_{k}$ (see Equation 2). (D) The PS encodings $\mathbf{X}_{r,k}$ are transformed using $f_{r,k}(\cdot)$, and initial cell features $\mathbf{Z}_{r}$ are embedded via $\varepsilon_{r}(\cdot)$, following Equations 3 and 4. (E) Rank-specific, neighborhood-aware representations $\mathbf{H}_{r}$ are learned using $\Theta_{r}(\cdot)$, as described in Equation 5. (F) The final task-specific readout $\varphi(\cdot)$ integrates all $\mathbf{H}_{r}$ representations (Equation 6).
@@
Let ${\mathcal{N}}_{C}$ be a collection of neighborhood functions. For each neighborhood ${\mathcal{N}}\in{\mathcal{N}}_{C}$, the Hasse graph decomposition of ${\mathcal{T}}$ enables the treatment of every ${\mathcal{J}}_{{\mathcal{N}}}$ as an independent component (see Figure 2A), each representing a unique set of connectivity relationships. $\mathbf{G}$ is a set of Positional and Structural Encoders (PSEs) such that
$\mathbf{G}\colon\{1,\dots,K\}\times\mathbf{J}_{\mathcal{C}}\longrightarrow\mathbb{R}^{N\times D_{k}},\quad\left(k,{\mathcal{J}}_{{\mathcal{N}}}\right)\longmapsto g_{k}\left({{\mathcal{J}}_{{\mathcal{N}}}\bigr{|}}_{k}\right).$ (1)
Here $g_{k}(\cdot):\mathbf{J}_{\mathcal{C}}\to\mathbb{R}^{|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}\times D_{k}}$ denotes a particular PSE, with $|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}$ being the number of nodes in the Hasse graph induced by neighborhood ${\mathcal{N}}$ corresponding to the target cells (illustrated in Figure 2B). Note that each $g_{k}(\cdot)$ maps to a different $D_{k}$, allowing PSEs to have different dimensions.
@@
The proposed neighborhood aggregation mechanism $\psi_{k}(\cdot)$ operates by grouping the encodings of cells with the same target rank $r$ to produce a unique embedding for each cell and each $k\in\{1,\dots,K\}$ (see Figure 2C). The embeddings for cells of rank $r$ are computed as:
$\mathbf{X}_{r,k}=\psi_{k}(\{g_{k}\left({{\mathcal{J}}_{{\mathcal{N}}}\bigr{|}}_{k}\right)\mid{\mathcal{N}}\in{\mathcal{N}}^{\to r}\}),$ (2)
where $\mathbf{X}_{r,k}\in\mathbb{R}^{N_{r}\times D_{k^{\prime}}}$, with $N_{r}$ denoting the number of $r$-cells and $D_{k^{\prime}}$ representing the output feature dimension for attribute $k$. The functions $\psi_{k}(\cdot)$ can also be interpreted as rank-level aggregators, where the aggregation operation can be any standard reduction method. Common choices include concatenation, sum, average, or even learnable transformations such as MLPs. The main advantage of using non-learnable aggregation functions for $\psi_{k}$ is that the embeddings $\mathbf{X}_{r,k}$ can be precomputed during dataset preprocessing, since they do not require learning.
Learning embeddings of cells.
In the HOPSE framework, each $r$-cell is associated with two complementary sources of information: (i) higher-order positional and structural encodings $\mathbf{X}_{r,k}$ aggregated from multiple neighborhoods, and (ii) initial features $\mathbf{Z}_{r}$ provided by the dataset or derived during the construction of the higher-order structure. Both types of features are independently projected into latent spaces through sets of learnable functions (see Figure 2D):
@@
where $f_{r,k}(\cdot)$ is applied to each feature $k$ of the $r$-cells, and $\varepsilon_{r}(\cdot)$ to their initial features.
@@
Next, to integrate the structural information with the cells’ initial attributes, the extracted and transformed higher-order PS encodings and the cell features are passed through the aggregator $\Theta_{r}(\cdot)$ (see Figure 2E) to obtain the final neighborhood-aware representations $\mathbf{H}_{r}\in\mathbb{R}^{N_{r}\times D_{r}}$:
@@
Since the structural information is now contained in the cell features, they can be treated independently. For example, for complex-level tasks where the network needs an output for the whole complex, the global transformation $\varphi(\cdot)$ combines the aggregated embeddings from all ranks (see Figure 2F). The overall output is
$\text{HOPSE}({\mathcal{T}})=\varphi\Bigl{(}\mathbf{H}_{0},\mathbf{H}_{1},\ldots,\mathbf{H}_{R}\Bigr{)}.$ (6)
4.2 Theoretical guarantees
The expressivity of GNNs is commonly analyzed through their correspondence with the Weisfeiler–Leman (WL) isomorphism test [26]. In the context of TDL, a generalization of WL over arbitrary neighborhoods of a combinatorial complex (CC) has been proposed, referred to as the CCWL test [39]. The following theorem establishes the expressivity of HOPSE with respect to it.
Theorem 4.1.
HOPSE is more powerful than the CCWL test in distinguishing non-isomorphic combinatorial complexes if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are injective for all $k\in{1,...,K}$ and for all $r\in{0,...,R}$.
The proof idea is that although HOPSE does not rely on message passing, with appropriately chosen neighborhoods, it can compute positional encodings using the same information as CCWL. Given the injectivity assumptions, this implies at least equal expressivity. The proof concludes by showing two CCs distinguished by HOPSE but not by CCWL. Full details are in Appendix C.
Another key property for models working on combinatorial data is cell permutation equivariance—the ability to yield consistent outputs under permutations of the cells. The following theorem establishes that HOPSE preserves this property under reasonable conditions:
Theorem 4.2.
HOPSE is cell permutation equivariant if the functions $\varepsilon_{r}(\cdot)$, $g_{k}(\cdot)$, $\psi_{k}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$ are cell permutation equivariant for all $k\in{1,\dots,K}$ and all $r\in{0,\dots,R}$.
This follows from the fact that compositions of permutation-equivariant functions remain equivariant. The proof is in Appendix C. While this requirement may seem strong, standard MLPs—though not equivariant to feature permutations—are equivariant to permutations of input sets, making them suitable functions for $\varepsilon_{r}(\cdot)$, $f_{r,k}(\cdot)$, and $\Theta_{r}(\cdot)$.
@@ 4.3 Realization of the framework
The practical implementation of the proposed framework requires defining a set of neighborhoods of interest ${\mathcal{N}}_{C}=\{{\mathcal{N}}_{1},\dots,{\mathcal{N}}_{T}\}$ (discussed in Section 5), a set of encoding functions $\mathbf{G}$ for deriving structural and positional features, neighborhood aggregation functions $\psi_{k}(\cdot)$, rank-specific processing functions $f_{r,k}(\cdot)$, and a structure mixing function $\Theta_{r}(\cdot)$. In particular, this work adopts $\psi_{k}(\cdot)=\mathrm{CONCAT}(\cdot)$ for aggregation, enabling each positional encoding to be handled independently. For both $f_{r,k}(\cdot)$ and $\Theta_{r}(\cdot)$, trainable multilayer perceptrons (MLPs) are used. The choice of encoding functions $\mathbf{G}$ defines the two proposed instantiations of the framework, HOPSE-M and HOPSE-G, offering a trade-off regarding the informativeness of extracted PS encoding and the computation cost to generate them. Additional implementation details are provided in Appendix D.
@@
This realization incorporates manually selected positional and structural encodings commonly used in graph learning. Specifically, $\mathbf{G}=\{g_{1},g_{2},g_{3},g_{4}\}$ includes: (i) eigenvector-based encodings, (ii) electrostatic potential-based encodings, (iii) random-walk-based encodings, and (iv) heat kernel signatures. These encodings are chosen to provide a fair comparison with HOPSE-G, as the GPSE model used in that variant was trained on the same set of positional and structural descriptors [13]. While informative, several of these encodings—especially those requiring eigendecomposition such as (i), (ii), and (iv)—are computationally intensive, with complexity ${\mathcal{O}}(n^{3})$ with $n$ the number of graph nodes.
@@
This variant leverages the pretrained GPSE network [13] to encode structural and positional information through a lightweight message-passing architecture. In this realization, $\mathbf{G}=\{g_{1}\}$ with $g_{1}=\textsc{GPSE}(\cdot)$. Unlike manual encodings that rely on computationally intensive operations such as eigendecomposition, GPSE produces latent descriptors using standard message passing. This makes HOPSE-G particularly suitable for large combinatorial complexes with many cells, offering a scalable alternative without significant loss in expressiveness.
@@
The presented implementation of HOPSE is designed for computational efficiency: no training is required up to and including Equation 2, allowing the matrices $\mathbf{X}_{r,k}$ to be fully precomputed in a preprocessing phase. By offloading all structure-dependent computations to preprocessing, HOPSE has complexity that scales linearly with the number of cells while retaining high expressive power, as established in Theorem 4.1. The efficiency of this design is further validated by experimental runtime results in Table 3. A comprehensive complexity analysis and a comparison with GCCN—a highly general higher-order architecture—are provided in Appendix E.
@@ 5 Experiments
This section empirically evaluates the proposed HOPSE framework, benchmarked against state-of-the-art TNNs. Both instantiations of the framework, HOPSE-M and HOPSE-G (see Section 4.3), are assessed on graph molecular datasets, as well as on expressivity and topological tasks. The standardized experimental protocol introduced in TopoBench [42] is adopted to ensure fair and reproducible comparisons through a comprehensive optimization and evaluation suite. The experimental setup is carefully designed to systematically address the following research questions:
@@
    Can the proposed HOPSE framework and its realizations (HOPSE-M and HOPSE-G) effectively approximate the performance of TNNs that leverage HOMP on standard TDL datasets?
@@
    Can HOPSE achieve competitive performance on synthetic datasets specifically constructed to evaluate the ability to capture and utilize topological properties?
@@
    How expressive HOPSE can get in practical scenarios compared with other GNNs and TNNs?
  4. Q4.
    To what extent does the reduced complexity of HOPSE, relative to HOMP-based methods, yield practical improvements in execution cost and scalability on standard TDL benchmarks?
@@ 5.1 Evaluation setup
First, the framework is evaluated on five standard benchmarking datasets commonly used in the TDL literature [42], comprising four graph classification tasks and one graph regression task. The classification tasks use datasets from the TUDataset collection [^3] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluationMUTAG, PROTEINS, NCI1, and NCI109 [35]. For regression, the ZINC dataset is used [27]. Following standard practice in TDL, structural lifting is applied to lift these graph datasets into higher-order domains—specifically, cycle-based lifting for the cell domain and clique complex lifting for the simplicial domain. Next, MANTRA family [2] represents purely topological datasets and tasks. It comprises over 43,000 triangulations of 2-manifolds. Among the multiple classification tasks that MANTRA datasets offer, this analysis focuses on the following three: predicting the homeomorphism name, predicting the orientability of a triangulated surface, and assigning a numerical label to the first three Betti numbers $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$. Lastly, the evaluation includes the CLS dataset [36], used to assess the expressivity of GNNs [16], which contains 150 graphs across 10 isomorphism classes. For further details, see Appendix F.
@@
The SaNN architecture is re-implemented due to unavailable source code [19] and included in all benchmarks. For the MANTRA family [2], baseline results are reported for the first time under the TopoBench protocol, ensuring consistency with the broader evaluation framework. Full details of the hyperparameter search and selected configurations are provided in Appendix F.2.
@@
To ensure fair comparisons, five stratified splits with a 50%/25%/25% train/validation/test ratio are considered for each dataset and task –except on ZINC, which uses predefined splits [27]. The optimal hyperparameter configuration is selected based on the average validation performance across five random seeds. Evaluation metrics include classification accuracy (MUTAG, PROTEINS, NCI1, NCI109), mean absolute error (MAE; ZINC), and F1 score (MANTRA). For each dataset, the mean and standard deviation are computed across the five test sets and reported in Table 1. Publicly available results are adopted from Telyatnikov et al. [42] and Papillon et al. [39] where applicable for graph datasets.

### v2

@@ 4.1 Methodology
This section describes HOPSE, a method for learning representations of (CCs) through the caracterization of it’s pieces. Characterizing a CC amounts to determining properties or invariants. These are shared among equivalnt CCs and allow distinguishing non-equivalent CCs from each toher, however similar they may be. The Hasse decomposition facilitates this characterization as a CC is determined by it’s Hasse decomposition. This decomposition results in a set of graphs that determines the initial object. For the theoretical motivation refer to Appendix˜B.
This work leverages the previous theoretical insights and the extensive research in graph characterization, for example through PSEs, to employ Hasse graphs as proxy objects to characterize a CC and model higher-order interactions. In the spirit of (Frasca et al., 2020), HOPSE partially decouples the structural from the functional representation. The structure is given by the Hasse decomposition while the functional representation is given by the attributes over the complex. Learning both structure and features in parallel circumvents the high cost of mixing features through neighborhood relationships, the pillar behind HOMP. The full pipeline is illustrated in Figure˜2, with components referenced using the corresponding labels (A–F) throughout this section.
Figure 2: HOPSE pipeline. (A) Considering the collection of neighborhood functions ${\mathcal{N}}_{C}=\{{\mathcal{A}}_{0,1},{\mathcal{I}}_{0\to 1},{\mathcal{I}}_{2\to 1},{\mathcal{I}}_{0\to 2}\}$, the input combinatorial complex $\mathcal{T}$ is decomposed into the corresponding Hasse graphs expansions $\{\mathcal{G}_{{\mathcal{N}}}\}_{{\mathcal{N}}_{C}}$. (B) PSEs are computed for each Hasse graph, $\mathbf{G}(\mathcal{G}_{\mathcal{N}_{i}})$, as defined in Equation˜1. (C) Higher-order PSEs are derived via the target rank-aware aggregation $\psi_{k}$ (see Equation˜2). (D) The PSEs $\mathbf{X}_{r,k}$ are transformed using $f_{r,k}(\cdot)$, and initial cell features $\mathbf{Z}_{r}$ are embedded via $\varepsilon_{r}(\cdot)$, following Equations˜3 and 4. (E) Rank-specific, structure-aware representations $\mathbf{H}_{r}$ are learned using $\Theta_{r}(\cdot)$, as described in Equation˜5. (F) The final task-specific readout $\varphi(\cdot)$ integrates all $\mathbf{H}_{r}$ representations (Equation˜6).
@@
Let ${\mathcal{N}}_{S}$ be a collection of neighborhood functions. For each neighborhood function ${\mathcal{N}}\in{\mathcal{N}}_{S}$, the Hasse graph decomposition of $\mathcal{T}$} enables the treatment of every ${\mathcal{G}}_{{\mathcal{N}}}$ as an independent component (see Figure˜2A), each representing a unique set of connectivity relationships. $\mathbf{G}$ is a set of Positional and Structural Encoders (PSEs) such that
$\mathbf{G}\colon\{1,\dots,K\}\times\mathcal{G}_{S}\longrightarrow\mathbb{R}^{N\times D_{k}},\quad\left(k,\mathcal{G}_{\mathcal{N}}\right)\longmapsto g_{k}\left(\mathcal{G}_{\mathcal{N}}\right).$ (1)
Here $g_{k}(\cdot):\mathcal{G}_{S}\to\mathbb{R}^{|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}\times D_{k}}$ denotes a particular PSE, with $|{\mathcal{C}}_{\mathcal{N}}|_{\text{target}}$ being the number of nodes in the Hasse graph constructed from the neighborhood function $\mathcal{N}$} corresponding to the target cells (illustrated in Figure˜2B). Note that each $g_{k}(\cdot)$ maps to a different $D_{k}$, allowing PSEs to have different dimensions.
@@
The function $\psi_{k}(\cdot)$ groups encodings of cells with the same target rank $r$ to produce a embedding that characterizes each cell and for each $k\in\{1,\dots,K\}$ (see Figure˜2C). The embeddings for cells of rank $r$ are computed as:
$\mathbf{X}_{r,k}=\psi_{k}(\{\!\!\{g_{k}(\mathcal{G}_{\mathcal{N}})\mid{\mathcal{N}}\in{\mathcal{N}}^{\to r}\}\!\!\}),$ (2)
where $\mathbf{X}_{r,k}\in\mathbb{R}^{N_{r}\times D_{k^{\prime}}}$, with $N_{r}$ denoting the number of $r$-cells and $D_{k^{\prime}}$ representing the output feature dimension for attribute $k$. The functions $\psi_{k}(\cdot)$ can also be interpreted as rank-level aggregators, where the aggregation operation can be any standard reduction method over a multiset  [^2] Common choices include concatenation, sum, average, or even learnable transformations such as MLPs.. The main advantage of using non-learnable aggregation functions for $\psi_{k}$ is to precompute the embeddings $\mathbf{X}_{r,k}$ during dataset preprocessing.
Learning structure and attributes.
In the HOPSE framework, each $r$-cell is associated with two complementary sources of information: (i) neighborhood-aware PSEs $\mathbf{X}_{r,k}$ , and (ii) initial features $\mathbf{Z}_{r}$ provided by the dataset or derived during the construction of the higher-order structure. Both types of features are projected in parallel into latent spaces through sets of learnable functions (see Figure˜2D):
@@
where $f_{r,k}(\cdot)$ is applied to each structural feature matrix $\mathbf{X}_{r,k}$ of each $r$-cell, and $\varepsilon_{r}(\cdot)$ to their initial features.
@@
Next, to integrate the structural information with the cells’ initial attributes, $\Theta_{r}(\cdot)$ aggregates their latent representations (see Figure˜2E) to obtain the final neighborhood-aware representations $\mathbf{H}_{r}\in\mathbb{R}^{N_{r}\times D_{r}}$:
@@
Since the structural information is now contained in the cell features, they can be treated independently. For example, for complex-level tasks where the network needs an output for the whole complex, the global transformation $\varphi(\cdot)$ combines the aggregated embeddings from all ranks (see Figure˜2F). The overall output is
$\text{HOPSE}({\mathcal{T}})=\varphi\Bigl(\mathbf{H}_{0},\ldots,\mathbf{H}_{R}\Bigr).$ (6)
@@ 4.2 Realization of the framework
The practical implementation of the proposed framework requires defining a set of $T$ neighborhood functions of interest ${\mathcal{N}}_{S}=\{{\mathcal{N}}_{1},\dots,{\mathcal{N}}_{T}\}$ (discussed in Section˜5), a set of encoding functions $\mathbf{G}$ for deriving structural and positional features, neighborhood aggregation functions $\psi_{k}(\cdot)$, rank-specific processing functions $f_{r,k}(\cdot)$, and a structure mixing function $\Theta_{r}(\cdot)$. In particular, this work adopts $\psi_{k}(\cdot)=\mathrm{CONCAT}(\cdot)$ for aggregation, enabling each positional encoding to be handled independently. The functions $\epsilon_{r}(\cdot)$, $f_{r,k}(\cdot)$ and $\Theta_{r}(\cdot)$ are modeled as multilayer perceptrons (MLPs). The choice of functions $\mathbf{G}$ defines two instantiations of the framework, HOPSE-M and HOPSE-G, offering a trade-off between representational power  [^3] Since HOPSE-G uses message passing to generate the embeddings of the Hasse graphs it is not completely message passing free.. and computation cost. Additional implementation details are provided in Appendix˜D.
@@
This realization incorporates manually selected positional and structural encodings commonly used in graph learning. Specifically, $\mathbf{G}=\{g_{1},g_{2},g_{3},g_{4}\}$ includes: (i) eigenvector-based encodings, (ii) electrostatic potential-based encodings, (iii) random-walk-based encodings, and (iv) heat kernel signatures. Section˜F.2 contains an in depth description of the choice of PSEs.
@@
This variant leverages the pretrained GPSE network (Cantürk et al., 2023) to encode structural and positional information. In this realization, $\mathbf{G}=\{g_{1}\}$ with $g_{1}=\textsc{GPSE}(\cdot)$. Unlike manual encodings that rely on computationally intensive operations such as eigendecomposition, GPSE produces latent descriptors using standard message passing. Domain specific tasks can thus benefit from models pretrained in that domain.
Choice of ${\mathcal{N}}_{S}$.
The choice of the set of neighborhood functions ${\mathcal{N}}_{S}$ used by HOPSE is a critical design choice in the HOPSE framework. Following the theoretical motivation (Appendix˜B) extensive ablation studies show the empirical performance (Appendix˜H) and runtime behavior (Appendix˜I) of the models.
@@
The HOPSE framework is efficient. No training is required up to and including Equation˜2, allowing the matrices $\mathbf{X}_{r,k}$ to be fully precomputed in a preprocessing phase. By decoupling the learning procedure of features and structure, HOPSE has complexity that scales linearly with the number of cells.
@@ 5 Experiments
This section empirically evaluates the HOPSE framework against state-of-the-art TNNs. Both instantiations of the framework, HOPSE-M and HOPSE-G (see Section 4.2), are assessed on graph molecular datasets and topological tasks. This work adopts the standardized experimental protocol introduced in TopoBench (Telyatnikov et al., 2024) to ensure fair and reproducible comparisons through a comprehensive optimization and evaluation suite. The experimental setup is carefully designed to systematically address the following research questions:
@@
    Is HOMP required for standard TDL evaluation benchmarks ?
@@
    Is HOMP a hard requirement for tasks related to topological properties?
@@
    Does the lower computational complexity result in a reduced training and inference time ?
@@ 5.1 Evaluation setup
First, the framework is evaluated on five standard benchmarking datasets commonly used in the TDL literature (Telyatnikov et al., 2024), comprising four graph classification tasks and one graph regression task. The classification tasks use datasets from the TUDataset collection [^4] Although these datasets are relatively dated, they remain widely adopted as a benchmarking suite in TDL research, offering consistent baselines for comparative evaluation.MUTAG, PROTEINS, NCI1, and NCI109 (Morris et al., 2020), and the regression is ZINC-12K (Irwin et al., 2012). Following standard practice in TDL, structural lifting is applied to transform these graph datasets into higher-order domains—specifically, cycle-based lifting for the cell domain and clique complex lifting for the simplicial domain. Next, the MANTRA set of datasets (Ballester et al., 2025) represents purely topological datasets and tasks. It comprises over 43,000 triangulations of 2-manifolds. Among the multiple classification tasks that MANTRA datasets offer, this analysis focuses on the following three: predicting the homeomorphism type (NAME), predicting the orientability of a triangulated surface (ORIENT), and assigning a numerical label to the first three Betti numbers $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$. For further discussion of the chosen benchmarks refer to Section˜F.2.
@@
The benchmark includes several models from the simplicial and cellular domains. In the simplicial domain: SCN (Yang et al., 2022), SCCNN (Yang and Isufi, 2023), and GCCN (Papillon et al., 2024). In the cellular domain: CCCN (Hajij et al., 2020), CWN (Bodnar et al., 2021), and GCCN. This work also provides a re-implementation of the SaNN architecture Gurugubelli and Chepuri (2024), as its source code was unavailable. Additionally, the first baseline results for the MANTRA family (Ballester et al., 2025) under the TopoBench protocol. Full details on the hyperparameter search and selected configurations are provided in Appendix F.3.
@@
Publicly available results are adopted from Telyatnikov et al. (2024) and Papillon et al. (2024) where applicable for graph datasets. While for the rest, the TopoBench evaluation protocol was applied, which consists of stratified splits with 50%/25%/25% train/validation/test ratios, except on ZINC-12K, which uses predefined splits (Irwin et al., 2012). The performance for each dataset is the result of the test performance of the highest validation, all averaged over 5 seeds. Across all tables, results are shown as mean $\pm$ standard deviation. The best results per domain are highlighted in bold, the best overall in red and those within one standard deviation of the best are in blue. The evaluation metrics are classification accuracy for MUTAG, PROTEINS, NCI1, NCI109, mean absolute error (MAE) for ZINC-12K, and F1 score for MANTRA. This standardized protocol ensures all methods are evaluated on an even playing field, which may lead to results that differ from some prior publications. Please refer to Section˜F.2 for further details. Additionally, Appendix˜K contains the full parameter budget of the best models.